8. Youre on the campus of SHU (Statistics Heave n U) which has a student You wish to find out the proportion of confidence interval withi u percentage points and no prior estimate is known. How many students population of 15ccs . students that like math. You want a gy should you sample?
Answer #1

8)

Solution :

Given that,

\hat p = 0.5

1 - \hat p = 1 - 0.5 = 0.5

margin of error = E = 4% = 0.04

At 94% confidence level the z is ,

\alpha = 1 - 94% = 1 - 0.94 = 0.06

\alpha / 2 = 0.06 / 2 = 0.03

Z\alpha/2 = Z0.03 = 1.88

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p)

= (1.88 * 0.04)2 * 0.5 * 0.5

= 552.25 = 553

Students = 553